OpenStudy (abbycross167):

Can someone please help me to answer this math question? Please? I will fan and give a medal... Please help me to simplify this Rational Expression - 5 2x+5 -------- - ------------- x+5 x^2+9x+20

OpenStudy (peachpi):

you need to get a common denominator. Start by factoring \(x^2+9x+20\)

OpenStudy (abbycross167):

(5 + x)(4 + x)

OpenStudy (peachpi):

right. so now you have to make the first fraction have the same denominator by multiplying both numerator and denominator by (x + 4). Like this \[\frac{ 5(x+4) }{ (x+5)(x+4) }-\frac{ 2x+5 }{ (x+5)(x+4) }\] Then you can combine it into 1 fraction like this \[\frac{ 5(x+4)-(2x+5) }{ (x+4)(x+5) }\] Simplify the top

OpenStudy (abbycross167):

ok

OpenStudy (abbycross167):

3 ---- (x+4) ? @peachpi

OpenStudy (peachpi):

yes

OpenStudy (abbycross167):

so that's the answer?

OpenStudy (abbycross167):

Thank you so much for the help!! I really appreciate it :)